A thin circular ring of mass M and radius R is rotating about its axis with a constant angular velocity ѡ. Four objects each of mass m are kept gently to the opposite ends of two perpendicular diameters of the ring.What will be the angular velocity of the ring?

A thin  circular ring of mass M and radius R is rotating about its axis with a constant angular velocity  ѡ. Four objects each of mass m are kept gently to the opposite ends of two perpendicular diameters of the ring.What will be the angular velocity of the ring?

Grade:11

1 Answers

Arun
5 years ago
Using conservation of angular momentum initial angular momentum=final angular momentum.
MR^2*w = (MR^2 + 2mR^2)w'
Mw/(M + 2m) = w'
w' is the new angular velocity